OpenStudy (anonymous):

What is the 24th term of the arithmetic sequence where a1 = 8 and a9 = 56 ?

OpenStudy (a_clan):

general formula, An = A1 + (n-1)d where A1 is the first term, 'n' is the no. of terms in the series and 'd' is the common difference. Acc to ques, A9= 8+ (9-1)d 56= 8+8d d=6 now, this info can be used to solve further